OpenStudy (anonymous):

Solve showing steps: 2s-5(-3+4s)=-39

OpenStudy (anonymous):

(-5*-3)+(-5*4s)???

OpenStudy (anonymous):

use the foil method.. http://images.tutorvista.com/cms/images/83/foil-calculator.PNG

OpenStudy (anonymous):

-6s+8s-20s+15=-39???

OpenStudy (imtiaz7):

6s+8s^2+15-20s= -39 8s^2-14s= -39-15 8s^2-14s= 54

OpenStudy (anonymous):

then simplify, @Imtiaz7 already did it.. I thought we are not allowed to post the answers..

OpenStudy (imtiaz7):

@eLg try to solve this by factoriaztion

OpenStudy (anonymous):

anyone get s=3 ?????

OpenStudy (imtiaz7):

or it can be solved by quadratic formula

OpenStudy (anonymous):